OpenStudy (anonymous):

b) Determine the upper sums and lower sums of \( \Large f(x) = x^{3}\) over \(\Large I = [0,1]\) for equidistant partitions.

OpenStudy (anonymous):

you mean compute the reimann sum?

OpenStudy (anonymous):

i am not sure but i have a) part of question and there was a sum .. induction to prove etc..

OpenStudy (anonymous):

should i post part a) too here ?

OpenStudy (anonymous):

divide the interval in to \(n\) equal parts of length \(\frac{1}{n}\) then put \(x_0=0\) \(x_1=\frac{1}{n}\), \(x_2=\frac{2}{n}\) and in general \(x_k=\frac{k}{n}\)

OpenStudy (anonymous):

you want \[\sum_{k=0}^n(\frac{k}{n})^3\times \frac{1}{n}\] or \[\frac{1}{n^4}\sum_{k=0}^nk^3\]

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

look up the formula for the sum of n cubes, which is the square of the formula for the sum of \(n\) namely \[\left(\frac{n(n+1)}{2}\right)^2\]

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

so you get \[\frac{1}{n^4}\times \left(\frac{n(n+1)}{2}\right)^2\]

OpenStudy (anonymous):

take the limit as \(n\to \infty\) and you will get \(\frac{1}{4}\) because both numerator and denominator are polynomials of degree 4 and the ratio of the leading coefficients is \(\frac{1}{4}\)

OpenStudy (anonymous):

ancient arab mathematicians new this formula, i think around 1100

OpenStudy (anonymous):

*knew

OpenStudy (anonymous):

wow you know even all math where the root is coming from and which time :)

OpenStudy (anonymous):

@satellite73 thank you very much ..

OpenStudy (anonymous):

yw you might want to put in more details, but i think the idea is all there good luck

OpenStudy (anonymous):

hmm i hope i get points with this solution :) it should be enough..

OpenStudy (vishweshshrimali5):

@satellite73 kindly help me please this question has got me stuck for more than a month http://openstudy.com/study?login#/updates/4ff65196e4b01c7be8c9590f